# An image-computable model of speeded decision-making

### v1

@@ Abstract
Evidence accumulation models (EAMs) are the dominant framework for modeling response time (RT) data from speeded decision-making tasks. While providing a good quantitative description of RT data in terms of abstract perceptual representations, EAMs do not explain how the visual system extracts these representations in the first place. To address this limitation, we introduce the visual accumulator model (VAM), in which convolutional neural network models of visual processing and traditional EAMs are jointly fitted to trial-level RTs and raw (pixel-space) visual stimuli from individual subjects. Models fitted to large-scale cognitive training data from a stylized flanker task captured individual differences in congruency effects, RTs, and accuracy. We find evidence that the selection of task-relevant information occurs through the orthogonalization of relevant and irrelevant representations, demonstrating how our framework can be used to relate visual representations to behavioral outputs. Together, our work provides a probabilistic framework for both constraining neural network models of vision with behavioral data and studying how the visual system extracts representations that guide decisions.
@@ Introduction
Decision-making under time pressure is deeply embedded within the activities of daily life. To study the cognitive and neural processes underlying decision-making, psychologists fit computational models to response time (RT) data gathered from relatively simple cognitive tasks. Evidence accumulation models (EAMs), such as the diffusion-decision model [59, 58] and linear ballistic accumulator model (LBA) [8], are the most successful and widely-used computational models of decision-making. In the EAM framework, sensory evidence is accumulated (possibly with noise) until reaching a threshold, at which point an overt response is generated. As such, decision-making is described in terms of a small number of interpretable parameters that capture basic cognitive processes. Empirically, EAMs have been shown to capture RT distributions from a variety of cognitive tasks at the individual subject level [58, 80, 60, 8].
Despite this empirical success and theoretical merit, the simple characterization of decision-making encapsulated by EAMs results in somewhat restrictive applications [17]. For example, EAMs do not typically provide a detailed specification of how stimuli are transformed into evidence or of the motor processes that physically generate a response. On the sensory side, the modeler instead converts the raw high-dimensional stimulus into a small number of abstract perceptual features that drive the rate of evidence accumulation. In the context of visual tasks, as we study here, this implies that EAMs are not amenable to studying the visual processing steps underlying decision-making, despite decades of experimental and theoretical vision research. It is also difficult to model tasks with complex (e.g. naturalistic) visual stimuli with EAMs, since the relevant visual features must be specified by hand.
Here, we integrate convolutional neural network (CNN) models for visual feature extraction with traditional accumulator models of speeded decision-making in a framework we call the visual accumulator model (VAM). We use CNNs since they are image-computable: they accept arbitrary images as input, such as visual stimuli from cognitive tasks. CNNs also capture key features of biological vision [41, 85, 37]. Each layer of the CNN applies a convolution, nonlinearity, and pooling operation to the output of the previous layer with a bank of trainable filters. In early network layers, this results in spatially-organized units with local receptive fields, analogous to the retinotopic organization of the early mammalian visual pathway. The concatenation of multiple layers forms a hierarchy in which progressively more abstract features are extracted in deeper layers, in a way that is globally similar to the primate visual cortical hierarchy.
Two features of the VAM distinguish our framework from other models in which neural networks were used to generate RTs or RT proxies [71, 77, 39, 55, 22]. First, the VAM generates RTs via a traditional EAM, a component that prior models lack. The VAM therefore inherits the advantages in interpretability afforded by EAMs and allows us to connect to the rich decision-making literature from psychology. Second, the CNN and EAM components of the VAM are jointly fitted to the RT, choice, and raw (pixel-space) visual stimuli from a single participant in a unified Bayesian framework. Thus, in contrast to previous models, the representations learned by the CNN and EAM parameters are directly constrained by behavioral data. We leverage these advantages to explore how abstract, task-relevant information is extracted from raw sensory inputs, and investigate how the behavioral phenomenon of congruency effects arises as a consequence of the representation geometry learned by the CNN. In doing so, our framework also addresses one of the main criticisms of deep neural network models of vision that are optimized to perform particular tasks (e.g. object identification): these models account for few results from psychology [5, 44, 18, 42, 2]. Indeed, as we will show, fitting the CNN with human data—rather than optimizing the model to perform a task—has significant consequences for the representations learned by the model.
@@
We first describe the task—Lost in Migration (LIM), available as part of the Lumosity cognitive training platform—which will serve to ground the discussion of the model. LIM is a stylized version of the well-known arrows flanker task used in the study of cognitive control and visual attention (Fig. 1A; [73]). In LIM, participants are shown stimuli composed of several arrow-shaped birds. The task is to indicate the direction of the central bird (the target) using the arrow keys on the keyboard, ignoring the direction of the surrounding birds (the flankers). Participants engage with the task at home and are rewarded via a composite score that takes into account both speed and accuracy.
@@
As with other flanker tasks, a given trial is said to be congruent if targets and flankers are oriented in the same direction, and incongruent otherwise. A consistent observation from studies employing the flanker task and related conflict tasks are congruency effects: participants are slower and less accurate on incongruent trials [16, 69, 74]. Congruency effects are considered to index a specific aspect of cognitive control: they indicate the extent to which an individual can selectively attend to task-relevant information and ignore irrelevant information.
The visual accumulator is an image-computable EAM, composed of a CNN and EAM chained together (Fig. 1B). Minimally processed (pixel-space) visual stimuli are provided as inputs to the CNN. The outputs of the CNN correspond to the mean rates at which evidence is accumulated (the drift rates), one for each possible response (four in the case of LIM). The EAM then generates choices and RTs through a noisy evidence accumulation process. For each participant in the dataset, we fit one such model (CNN + EAM) jointly using that participant’s visual stimuli, RTs, and choices. Building on prior work [38, 11], we developed an automatic differentiation variational inference (ADVI) algorithm that simultaneously optimizes the parameters of the CNN and learns the posterior distribution over the LBA parameters (Methods).
The particular EAM we adopt is the linear ballistic accumulator model (LBA [8]; Fig. 1B) since it can be applied to tasks with more than two possible responses, though the general VAM framework is compatible with other EAMs that have a closed-form or easily approximated likelihood. The LBA parameters fitted in the VAM are the decision threshold $b$, the non-decision time $t_{0}$, and a parameter $A$ that controls the dispersion of the initial accumulator values (the drift rate means are fitted implicitly via the CNN). We used a seven layer CNN architecture (6 convolutional layers, 1 fully-connected layer) in the VAM (Fig. 1B). To speed up the training process, the first two convolutional layers were initialized with the parameters from a 16-layer VGG CNN trained to classify the ImageNet dataset [70, 12].
@@ Results
We also examined whether the fitted models captured demographic effects, focusing on the well-established slowing of RTs that occurs with age [50, 23, 57]. Mean RTs began increasing around age 50, effects captured by the fitted models (Fig. 2G). Consistent with prior work in a variety of decision-making tasks [72, 67, 57, 20], we found that an age-dependent increase in the non-decision time ($t_{0}$) component of the response contributed to longer RTs in the models from older adults (Fig. S2A). In contrast to these prior studies, we did not observe increased response caution (measured as $b-A$) in the models from older adults (Fig. S2B). This discrepancy could be explained by differences in the task design or the fact that our participants had practiced the task substantially more than in typical studies [72]. We also observed an age-dependent reduction in the target drift rates and no age-dependence of the flanker drift rates (Fig. S2C–D), findings that have received mixed support in the literature [72, 67, 57, 20, 3].
@@
While the analyses above demonstrate that the mean congruency effects from the participants were captured well by the VAMs, we observed a mismatch between model and participant behavior for other commonly used behavioral indices that take into account additional information from the RT distribution: RT delta plots and conditional accuracy functions [83, 30, 63, 78, 82] (Fig. S4A–B). However, when we fitted a separate model to trials from each of $n$=5 RT quantiles from a given participant’s data, quantitative fits improved and qualitative trends were captured for both delta plots and conditional accuracy functions (Fig. S4C–D, Methods). Thus, a small modification of the VAM training paradigm can be used to investigate how congruency effects change as a function of RT. However, a thorough investigation of these effects is outside the scope of this manuscript, and we focus on the models fit to all trials in the remainder of our study.
@@
To perform LIM well, the models must learn representations that select the task-relevant information (target direction) and diminish the influence of irrelevant information (flanker direction, stimulus layout, and stimulus position). To investigate these representations, we analyzed a set of stimulus features in each CNN layer, derived from the models’ responses to a holdout set of LIM stimuli (separate from the training stimuli; Fig. 3A, Methods). The features in each layer form an $N\times K_{l}$ matrix, where $N$ is the number of stimuli in the holdout set (5000) and $K_{l}$ is the number of active units in layer $l$ (a variable fraction of units in each layer did not respond to any stimuli and were excluded from the feature matrix).
@@
We characterized the emergence of target selectivity in the model by quantifying how well the target direction could be decoded from the feature matrix of a given layer with a linear support vector machine [66] (SVM; Fig. 3B). Note that only incongruent trials were used in these analyses, since the classifier could use the flanker direction to classify the target on congruent trials, artificially inflating performance. Decoding performance on holdout data for target direction was near chance (27%) in the first network layer and increased to nearly perfect decoding ($\geq$ 97%) at layer 4, with a sharp increase between the second and third convolutional layers (Fig. 3B). The increase in target decoding accuracy from shallower to deeper network layers is generally consistent with neural recording studies in mammals and other neural networks studies that document more accurate decoding of abstract variables (e.g. object or category identity) in higher visual/cortical regions and deeper neural network layers [47, 76, 7].
@@
We also investigated target selectivity at the single-unit level by quantifying the mutual information between each unit’s activity and target direction [76, 47] (Fig. 3C). In contrast to the population-level decoding results, information for target direction exhibited a non-monotonic "hunchback" profile across the convolutional layers, then increased sharply in the final fully-connected layer. The observation that single-unit information for target direction decreased between the fourth and final convolutional layers while population-level decoding remained high is especially noteworthy in that it implies a transition from representing target direction with specialized "target neurons" to a more distributed, ensemble-level code. Notably, a similar transition in coding properties takes place along the cortical hierarchy, with higher-order cortical regions exhibiting a reduction in units with "pure" selectivity and a corresponding increase in units with mixed selectivity for multiple task or stimulus features [64, 45, 46]. The high proportion of units with mixed selectivity results in a high-dimensional representation in which all task-relevant information can easily be extracted with simple linear decoders [10, 51].
@@
Consistent with these ideas, we found that the dimensionality of target representations as measured by the participation ratio (Methods) increased sharply in the last two convolutional layers (Conv5–Conv6), paralleling the reduction in single-unit information for target direction in these layers (Fig. 3D). The high-dimensional representation observed in these layers may enable the VAM to capture the rich stimulus-behavior dependencies present in the participant data. The reduction in dimensionality in the final fully-connected layer, and concomitant increase in single-unit information for target direction, may reflect a strong constraint to select the correct target direction imposed by the task. Notably, the hunchback-shaped profile we observed in the dimensionality of representations has also been observed along visual cortical regions of the rat ventral stream and in other neural network studies [1, 47].
@@
Figure 3: Neural representations of target direction. A) Schematic of the CNN features extracted from each network layer. B) Decoding accuracy of stimulus target direction. C) Normalized mutual information for target direction conveyed by single units, averaged across units. Mutual information was normalized by the entropy of the target direction distribution. D) Dimensionality of target representations as measured by the participation ratio of the target-centered feature covariance matrix. E) Proportion of units exhibiting selectivity for target direction. Panels B–E show the average across $n$ = 75 models; error bars correspond to bootstrap 95% confidence intervals.
@@
How do the models’ visual representations enable target selectivity for stimuli that vary along several irrelevant dimensions? The mammalian visual system solves the analogous problem of visual object recognition through representations that become increasingly invariant to different object positions, sizes, and lighting conditions [66, 14]. To determine whether the models learned representations that share these characteristics, we initially assessed whether the model representations for target direction are indeed invariant (or more generally, tolerant), to variation in flanker direction, stimulus layout, and horizontal/vertical position. Alternatively, the models could have learned a coding scheme in which different representations are responsible for selecting the target in each distracter context, e.g. with units that respond to the conjunction of particular target direction and layout combinations.
To assess the degree of representation tolerance, we quantified how well a linear SVM trained to classify target direction in one distracter context generalized to a different distracter context, where the distracter context is defined by a particular type of task-irrelevant information [66, 4]. For example, to assess the degree of tolerance to flanker direction, we split trials into a training set where the flanker direction was fixed to a particular value (e.g. down), and a generalization set with all other flanker directions (e.g. flanker direction = left/right/up). The performance of the classifier on the generalization set measures the extent to which the target representations are tolerant (or invariant) to variability in a given type of distracting information.
@@
We also examined suppression at the single-unit level by quantifying the mutual information between each unit’s activations and the values of each task-irrelevant stimulus feature [76] (Fig. 4C). The mutual information for a given stimulus feature was normalized by the entropy of the feature distribution to facilitate comparisons between the features [47]. In agreement with the population-level decoding analyses, information for the irrelevant stimulus features exhibited a pronounced hunchback profile in the progression from shallower to deeper network layers. It is noteworthy that the suppression of irrelevant information is more pronounced at the single-unit level in that decoding accuracy remains relatively high in later layers, particularly for flanker direction. This parallels the findings discussed above for target direction, and again suggests a transition from a simple code with populations of units that are selective for particular stimulus features to a more distributed, ensemble-level code.
Figure 4: Suppression of task-irrelevant information and tolerance in task-relevant representations. A) Decoding accuracy of stimulus target direction in a new distracter context (generalization performance). Context was defined by the values of a given stimulus feature (flanker direction, layout, horizontal/vertical position). B) Decoding accuracy of irrelevant stimulus features. C) Normalized mutual information for irrelevant stimulus features conveyed by single units, averaged across units. For each stimulus feature, the mutual information was normalized by the entropy of the stimulus feature distribution. All panels show the average across $n$ = 75 models; error bars correspond to bootstrap 95% confidence intervals.
@@
A noteworthy feature of visual attention is that the selectivity and tolerance identified above is not absolute: irrelevant information cannot be filtered out completely, as illustrated by the congruency effects observed in the flanker task and related conflict tasks. A common framework for understanding these behavioral phenomena posits that task-relevant and irrelevant information compete for control over response execution, and that congruency effects arise from incomplete suppression of irrelevant information [78, 83]. Motivated by these theories, we investigated whether the degree of suppression of irrelevant information in the trained models was correlated with congruency effects across the models we analyzed.
@@
Given this somewhat surprising negative result, we were motivated to consider an alternative account of congruency effects, one that takes into account the relative geometry of the task-irrelevant (flanker) and task-relevant (target) information. In particular, we considered the possibility that task-relevant and irrelevant information could be orthogonalized in the high-dimensional space of neural activity, such that task-relevant information is shielded from distracter interference. The general idea that neural representations for different types of information can be orthogonalized to prevent interference has received support from a number of neural recording studies [40, 52, 19, 32, 65].
To examine whether target and flanker representations are orthogonalized, we first defined target and flanker subspaces from the target direction and flanker direction classifiers used in the decoding analyses described above (Methods). The target direction classifier for a given network layer implicitly defines four decoding vectors, one for each target direction [4, 40]. We define the subspace of the $K_{l}$-dimensional feature space spanned by these four vectors as the target subspace; the flanker subspace is defined analogously.
To measure the orthogonality between target and flanker subspaces, we calculated the average of the cosine of the principal angles between the target and flanker subspaces, a metric we refer to as subspace alignment (Methods). Principal angles generalizes the idea of angles between lines or planes to arbitrary dimensions [88, 31]. The subspace alignment metric has a simple interpretation: it is equal to one if the subspaces are completely parallel and zero if they are completely orthogonal.
@@
Researchers who use neural network models to study neural representations typically optimize the model to perform a task, rather than fit the model to behavioral data from the task as we do here [87, 86, 81, 45], cf. [29, 13, 75]. This raises the possibility that these two training paradigms induce different representations in the models. To investigate this, we trained CNNs to perform LIM (i.e. output the direction of the target) by minimizing the standard cross-entropy loss function used in image classification tasks. Other than the training objective, all aspects of the optimization algorithm, CNN architecture, initialization, and training data were the same as those used to train the VAM. We trained one task-optimized model for each VAM, using data from the same participants ($n$ = 75 task-optimized models).

### v2

@@ Abstract
Evidence accumulation models (EAMs) are the dominant framework for modeling response time (RT) data from speeded decision-making tasks. While providing a good quantitative description of RT data in terms of abstract perceptual representations, EAMs do not explain how the visual system extracts these representations in the first place. To address this limitation, we introduce the visual accumulator model (VAM), in which convolutional neural network models of visual processing and traditional EAMs are jointly fitted to trial-level RTs and raw (pixel-space) visual stimuli from individual subjects in a unified Bayesian framework. Models fitted to large-scale cognitive training data from a stylized flanker task captured individual differences in congruency effects, RTs, and accuracy. We find evidence that the selection of task-relevant information occurs through the orthogonalization of relevant and irrelevant representations, demonstrating how our framework can be used to relate visual representations to behavioral outputs. Together, our work provides a probabilistic framework for both constraining neural network models of vision with behavioral data and studying how the visual system extracts representations that guide decisions.
@@ Introduction
Decision-making under time pressure is deeply embedded within the activities of daily life. To study the cognitive and neural processes underlying decision-making, psychologists fit computational models to response time (RT) data gathered from relatively simple cognitive tasks. Evidence accumulation models (EAMs), such as the diffusion-decision model [ratcliff2008, ratcliff1978memory] and linear ballistic accumulator model (LBA) [brown2008], are the most successful and widely-used computational models of joint RT and choice data from decision-making tasks. In the EAM framework, sensory evidence is accumulated (possibly with noise) until reaching a threshold, at which point an overt response is generated. As such, decision-making is described in terms of a small number of interpretable parameters that capture basic cognitive processes. Empirically, EAMs have been shown to capture RT distributions from a variety of cognitive tasks at the individual subject level [ratcliff1978memory, usher2001, ratcliff19982choice, brown2008].
Despite this empirical success and theoretical merit, the simple characterization of decision-making encapsulated by EAMs results in somewhat restrictive applications [evans2020]. In the context of visual tasks, EAMs do not typically provide a detailed specification of how raw visual stimuli are transformed into evidence and consequently do not explain the visual processing steps underlying decision-making. A number of recent modeling efforts have addressed this limitation by adapting convolutional neural network (CNN) models used in image classification tasks to the novel purpose of generating RTs or RT proxies [spoerer2020recurrentNNRTs, taylor2021NNRTs, kumbhar2020anytimepred, rafiei2024RTNet, goetschalckx2023RTrCNN, annis2021CNNLBA, holmes2020CNNDDM, trueblood2021CNNDDM], a strategy we also pursue here. CNNs are useful in this regard since they are image-computable—they accept arbitrary images as input—and capture important characteristics of biological vision [lindsay2021CNNreview, yamins2016perspective, kriegeskorte2015review]. Early CNN layers exhibit spatially-organized units with local receptive fields, analogous to the retinotopic organization of the early mammalian visual pathway. The concatenation of multiple layers forms a hierarchy in which progressively more abstract features are extracted in deeper layers, in a way that is globally similar to the primate visual cortical hierarchy.
Here, we integrate CNN models for visual feature extraction with traditional EAMs of speeded decision-making tasks in a framework we call the visual accumulator model (VAM). As in the prior models referenced above, the VAM accepts raw (pixel-space) visual stimuli as inputs and generates RTs and choices as outputs. The key feature of the VAM that distinguishes it from prior models is that the CNN and EAM parameters are jointly fitted to the RT, choice, and visual stimulus data from individual participants in a unified Bayesian framework. Thus, both the visual representations learned by the CNN and the EAM parameters are directly constrained by behavioral data. In contrast, prior models first optimize the CNN to perform the behavioral task, then separately fit a minimal set of high-level CNN parameters [rafiei2024RTNet] and/or the EAM parameters to behavioral data [annis2021CNNLBA, holmes2020CNNDDM, trueblood2021CNNDDM]. As we will show, fitting the CNN with human data—rather than optimizing the model to perform a task—has significant consequences for the representations learned by the model.
We leverage the VAM to explore how abstract, task-relevant information is extracted from raw sensory inputs, and we investigate how the behavioral phenomenon of congruency effects arises as a consequence of the representation geometry learned by the CNN. In doing so, our framework also addresses one of the main criticisms of deep neural network models of vision that are optimized to perform particular tasks (e.g., object identification): these models account for few results from psychology [bowers2022, malhotra2022cnnblindness, fel2022harmonizing, linsley2017visfts, baker2018cnn_vis].
@@
We first describe the task—Lost in Migration (LIM), available as part of the Lumosity cognitive training platform—which will serve to ground the discussion of the model. LIM is a stylized version of the well-known arrows flanker task used in the study of cognitive control and visual attention (Fig. 1A; [stoffels1988flanker]). In LIM, participants are shown stimuli composed of several arrow-shaped birds. The task is to indicate the direction of the central bird (the target) using the arrow keys on the keyboard, ignoring the direction of the surrounding birds (the flankers). Participants engage with the task at home and are rewarded via a composite score that takes into account both speed and accuracy.
@@
As with other flanker tasks, a given trial is said to be congruent if targets and flankers are oriented in the same direction, and incongruent otherwise. A consistent observation from studies employing the flanker task and related conflict tasks are congruency effects: participants are slower and less accurate on incongruent trials [eriksen1974flanker, simon1982, stroop1935]. Congruency effects are considered to index a specific aspect of cognitive control: they indicate the extent to which an individual can selectively attend to task-relevant information and ignore irrelevant information.
The visual accumulator is an image-computable EAM, composed of a CNN and EAM chained together (Fig. 1B). Minimally processed (pixel-space) visual stimuli are provided as inputs to the CNN. The outputs of the CNN correspond to the mean rates at which evidence is accumulated (the drift rates), one for each possible response (four in the case of LIM). The EAM then generates choices and RTs through a noisy evidence accumulation process. For each participant in the dataset, we fit one such model (CNN + EAM) jointly using that participant’s visual stimuli, RTs, and choices. Building on prior work [kucukelbir2017, dao2022], we developed an automatic differentiation variational inference (ADVI) algorithm that simultaneously optimizes the parameters of the CNN and learns the posterior distribution over the LBA parameters (Methods).
The particular EAM we adopt is the linear ballistic accumulator model (LBA [brown2008]; Fig. 1B) since it can be applied to tasks with more than two possible responses, though the general VAM framework is compatible with other EAMs that have a closed-form or easily approximated likelihood. The LBA parameters fitted in the VAM are the decision threshold $b$, the non-decision time $t_{0}$, and a parameter $A$ that controls the dispersion of the initial accumulator values (the drift rate means are fitted implicitly via the CNN). We used a seven layer CNN architecture (6 convolutional layers, 1 fully-connected layer) in the VAM (Fig. 1B). To speed up the training process, the first two convolutional layers were initialized with the parameters from a 16-layer VGG CNN trained to classify the ImageNet dataset [simonyan2015, deng2009].
@@ Results
We also examined whether the fitted models captured demographic effects, focusing on the well-established slowing of RTs that occurs with age [nettelbeck1992speed, gottsdanker1982age, ratcliff2001ageDDM]. Mean RTs began increasing around age 50, effects captured by the fitted models (Fig. 2G). Consistent with prior work in a variety of decision-making tasks [steyvers2019ebbflow, servant2020ageDMC, ratcliff2001ageDDM, forstmann2011ageLBA], we found that an age-dependent increase in the non-decision time ($t_{0}$) component of the response contributed to longer RTs in the models from older adults (Fig. S2A). In contrast to these prior studies, we did not observe increased response caution (measured as $b-A$) in the models from older adults (Fig. S2B). This discrepancy could be explained by differences in the task design or the fact that our participants had practiced the task substantially more than in typical studies [steyvers2019ebbflow]. We also observed an age-dependent reduction in the target drift rates and no age-dependence of the flanker drift rates (Fig. S2C–D), findings that have received mixed support in the literature [steyvers2019ebbflow, servant2020ageDMC, ratcliff2001ageDDM, forstmann2011ageLBA, ben-david2014ageLBA].
@@
We also examined whether the VAMs captured two other commonly used behavioral metrics that take into account additional information from the RT distribution: RT delta plots and conditional accuracy functions [wildenberg2010inhibition, jong1994deltaplot, ridderinkhof2002suppression_article, ulrich2015DMC, white2011]. The participant RT delta plots showed that the congruency effect increased with longer RTs [pratte2021], a trend that was captured by the fitted VAMs (Fig. S4A). In contrast, we observed a mismatch between model and participant behavior for the conditional accuracy functions on incongruent trials (Fig. S4B). In particular, the participants tended to be less accurate on faster trials, while the models exhibited the opposite trend. This discrepancy may be explained by the lack of dynamic visual processing in the VAM, which has been proposed to account for the preponderance of errors on faster trials often observed in conflict tasks [white2011, wildenberg2010inhibition]. In the remainder of our analyses, we focus on the mean congruency effect, leaving a full accounting of such dynamic error patterns for future modeling work.
@@
To perform LIM well, the models must learn representations that select the task-relevant information (target direction) and diminish the influence of irrelevant information (flanker direction, stimulus layout, and stimulus position). To investigate these representations, we presented the model with a holdout set of LIM stimuli (separate from the training stimuli) and analyzed the resulting unit activity in each CNN layer [hohman2020summit, zeiler2013featurevis] (Fig. 3A, Methods). The activations from a given layer $l$ form an $N\times K_{l}$ matrix, where $N$ is the number of stimuli in the holdout set (5000) and $K_{l}$ is the number of active units in layer $l$ (a variable fraction of units in each layer did not respond to any stimuli and were excluded from the activation matrix).
@@
To characterize the emergence of target selectivity, we adopted analysis techniques from the neuroscience literature that aim to determine what stimulus properties are coded by population-level neural activity, analogous to the high-dimensional CNN activations studied here. Specifically, we quantified how well the target direction could be decoded from the activation matrix of a given layer with a linear support vector machine [rust2010, koren2020, hung2005IT] (SVM; Fig. 3B). Note that only incongruent trials were used in these analyses, since the classifier could use the flanker direction to classify the target on congruent trials, artificially inflating performance. Decoding performance on holdout data for target direction was near chance (27%) in the first network layer and increased to nearly perfect decoding ($\geq$ 97%) at layer 4, with a sharp increase between the second and third convolutional layers (Fig. 3B). The increase in target decoding accuracy from shallower to deeper network layers is generally consistent with neural recording studies in mammals and other neural networks studies that document more accurate decoding of abstract variables (e.g., object or category identity) in higher visual/cortical regions and deeper neural network layers [muratore2022prune, tafazoli2017tolerance, brincat2018taskrelatedneural].
@@
We also investigated target selectivity at the single-unit level by quantifying the mutual information between each unit’s activity and target direction [tafazoli2017tolerance, muratore2022prune] (Fig. 3C). In contrast to the population-level decoding results, information for target direction exhibited a non-monotonic "hunchback" profile across the convolutional layers, then increased sharply in the final fully-connected layer. The observation that single-unit information for target direction decreased between the fourth and final convolutional layers indicates that the units become progressively less selective for particular target directions. Since population-level decoding remained high in these layers, this suggests a transition from representing target direction with specialized "target neurons" to a more distributed, ensemble-level code. Notably, a similar transition in coding properties takes place along the cortical hierarchy, with higher-order cortical regions exhibiting a reduction in units with "pure" selectivity and a corresponding increase in units with mixed selectivity for multiple task or stimulus features [rigotti2013selectivity, mante2013, meister2013multiplex]. The high proportion of units with mixed selectivity results in a high-dimensional representation in which all task-relevant information can easily be extracted with simple linear decoders [cover1965, pagan2013untangling].
@@
Consistent with these ideas, we found that the dimensionality of target representations as measured by the participation ratio (Methods) increased sharply in the last two convolutional layers (Conv5–Conv6), paralleling the reduction in single-unit information for target direction in these layers (Fig. 3D). The high-dimensional representation observed in these layers may enable the VAM to capture the rich stimulus-behavior dependencies present in the participant data. The reduction in dimensionality in the final fully-connected layer, and concomitant increase in single-unit information for target direction, may reflect a strong constraint to select the correct target direction imposed by the task. Notably, the hunchback-shaped profile we observed in the dimensionality of representations has also been observed along visual cortical regions of the rat ventral stream and in other neural network studies [ansuini2019intrinsic, muratore2022prune].
@@
Figure 3: Neural representations of target direction. A) Schematic of the CNN activations extracted from each network layer. Each layer yields a $N\times K_{l}$ activation matrix, where $N$ is the number of stimuli and $K_{l}$ is the number of active units (i.e., feature dimensions) in layer $l$. B) Decoding accuracy of stimulus target direction. C) Normalized mutual information for target direction conveyed by single units, averaged across units. Mutual information was normalized by the entropy of the target direction distribution (possible range = [0, 1]). D) Dimensionality of target representations as measured by the participation ratio of the target-centered activation covariance matrix. E) Proportion of units exhibiting selectivity for target direction. Panels B–E show the average across $n$ = 75 models; error bars correspond to bootstrap 95% confidence intervals.
@@
How do the models’ visual representations enable target selectivity for stimuli that vary along several irrelevant dimensions? The mammalian visual system solves the analogous problem of visual object recognition through representations that become increasingly invariant to different object positions, sizes, and lighting conditions [rust2010, dicarlo2012objectrecog]. To determine whether the models learned representations that share these characteristics, we initially assessed whether the model representations for target direction are indeed invariant (or more generally, tolerant), to variation in flanker direction, stimulus layout, and horizontal/vertical position. Alternatively, the models could have learned a coding scheme in which different representations are responsible for selecting the target in each distracter context, e.g., with units that respond to the conjunction of particular target direction and layout combinations.
To assess the degree of representation tolerance, we quantified how well a linear SVM trained to classify target direction in one distracter context generalized to a different distracter context, where the distracter context is defined by a particular type of task-irrelevant information [rust2010, bernardi2020abstraction]. For example, to assess the degree of tolerance to flanker direction, we split trials into a training set where the flanker direction was fixed to a particular value (e.g. down), and a generalization set with all other flanker directions (e.g., flanker direction = left/right/up). The performance of the classifier on the generalization set measures the extent to which the target representations are tolerant (or invariant) to variability in a given type of distracting information.
@@
We also examined suppression at the single-unit level by quantifying the mutual information between each unit’s activations and the values of each task-irrelevant stimulus feature [tafazoli2017tolerance] (Fig. 4C). The mutual information for a given stimulus feature was normalized by the entropy of the feature distribution to facilitate comparisons between the different stimulus features [muratore2022prune]. In agreement with the population-level decoding analyses, information for the irrelevant stimulus features exhibited a pronounced hunchback profile in the progression from shallower to deeper network layers. It is noteworthy that the suppression of irrelevant information is more pronounced at the single-unit level in that decoding accuracy remains relatively high in later layers, particularly for flanker direction. This parallels the findings discussed above for target direction, and again suggests a transition from a simple code with populations of units that are selective for particular stimulus features to a more distributed, ensemble-level code.
Figure 4: Suppression of task-irrelevant information and tolerance in task-relevant representations. A) Decoding accuracy of stimulus target direction in a new distracter context (generalization performance). Context was defined by the values of a given stimulus feature (flanker direction, layout, horizontal/vertical position). B) Decoding accuracy of irrelevant stimulus features. C) Normalized mutual information for irrelevant stimulus features conveyed by single units, averaged across units. For each stimulus feature, the mutual information was normalized by the entropy of the stimulus feature distribution (possible range = [0, 1]). All panels show the average across $n$ = 75 models; error bars correspond to bootstrap 95% confidence intervals.
@@
A noteworthy feature of visual attention is that the selectivity and tolerance identified above is not absolute: irrelevant information cannot be filtered out completely, as illustrated by the congruency effects observed in the flanker task and related conflict tasks. A common framework for understanding these behavioral phenomena posits that task-relevant and irrelevant information compete for control over response execution, and that congruency effects arise from incomplete suppression of irrelevant information [ulrich2015DMC, wildenberg2010inhibition]. Motivated by these theories, we investigated whether the degree of suppression of irrelevant information in the trained models was correlated with congruency effects across the models we analyzed.
@@
Given this somewhat surprising negative result, we were motivated to consider an alternative account of congruency effects, one that takes into account the relative geometry of the task-irrelevant (flanker) and task-relevant (target) information. In particular, we considered the possibility that task-relevant and irrelevant information could be orthogonalized in the high-dimensional space of neural activity, such that task-relevant information is shielded from distracter interference. The general idea that neural representations for different types of information can be orthogonalized to prevent interference has received support from a number of neural recording studies [libby2021orthogonal, panichello2021WM, flesch2022orthogonal, kaufman2014nullspace, ritz2024orthogonal].
To examine whether target and flanker representations are orthogonalized, we first defined target and flanker subspaces from the target direction and flanker direction classifiers used in the decoding analyses described above (Methods). The target direction classifier for a given network layer implicitly defines four decoding vectors, one for each target direction [bernardi2020abstraction, libby2021orthogonal]. We define the subspace of the $K_{l}$-dimensional feature space spanned by these four vectors as the target subspace; the flanker subspace is defined analogously.
To measure the orthogonality between target and flanker subspaces, we calculated the average of the cosine of the principal angles between the target and flanker subspaces, a metric we refer to as subspace alignment (Methods). Principal angles generalizes the idea of angles between lines or planes to arbitrary dimensions [zhu2013principal_angles, jordan1875principal_angles]. The subspace alignment metric has a simple interpretation: it is equal to one if the subspaces are completely parallel and zero if they are completely orthogonal.
@@
Researchers who use neural network models to study neural representations typically optimize the model to perform a task, rather than fit the model to behavioral data from the task as we do here [yang2019, yamins2014, wang2018, mante2013], cf. [jaffe2023taskdyva, dezfouli2019, sussillo2015]. This raises the possibility that these two training paradigms induce different representations in the models. To investigate this, we trained CNNs to perform LIM (i.e., output the direction of the target) by minimizing the standard cross-entropy loss function used in image classification tasks, where the training labels are given by the true direction of the target bird in each stimulus. The task-optimized CNNs were identical to those used in the VAMs, except that the outputs of the last layer were converted to softmax-scored probabilities for each direction rather than drift rates. Otherwise, all aspects of the optimization algorithm, CNN architecture, initialization, and training data for the task-optimized models were the same as those used to train the VAMs. We trained one task-optimized model for each VAM using stimulus data from the same participants ($n$ = 75 task-optimized models).